Circular No. 6206

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 Q1 A new comet—his seventeenth—has been discovered by William A.

Bradfield, Dernancourt, near Adelaide. Available observations:

1995 UT α_{2000} δ_{2000} m_{1} Observer
Aug. 17.410 11^{h}00.^{m}6 −14^{o}08′ 6  Bradfield
17.99  11 01.5 −13 27  5.5 Pizarro
18.28  11 03.5 −13 12  5  Drummond

W. Bradfield (Dernancourt). Tail > 1^{o}. Communicated by A. Beresford. G. Pizarro (European Southern Observatory). 1.0-m Schmidt. Rough po-

sition by R. M. West. 7′ coma, two wavy tails in p.a. 145^{o}, length 3^{o}. P. Bouchet, S. Benetti and others confirmed the comet, which was visible with the naked eye; 2^{o} tail.

J. Drummond (Gisborne, New Zealand). Communicated by W. Orchiston.

SUPERNOVA 1995W IN NGC 7650 A. Williams and R. Martin, Perth Observatory, report the discovery

on Aug. 5.65 UT by the Perth Astronomy Research Group’s automated supernova search of an apparent supernova (mag ∼ 16) located 9′′ west and 22′′ south of the center of NGC 7650 (α = 23^{h}25^{m}.3, δ = −57^{o}48′, equinox 2000.0). J. Greenhill and K. M. Hill, Canopus Observatory, University of Tasmania, write: “Instrumental magnitudes measured with the Canopus Observatory 1-m telescope on Aug. 10.6 UT show that the object was 1.0 mag fainter in v and much hotter (b − v = −0.26) than the closest bright star, located ∼ 1′.6 south and west of the supernova.”

S. Benetti and T. Augusteijn, European Southern Observatory; and A. Sarajedini, Kitt Peak National Observatory, communicate: “Inspection of a preliminarily-reduced CCD spectrum (range 396–793 nm, resolution 1 nm) obtained on Aug. 17.29 UT with the Danish 1.54-m telescope (+ DFOSC) at La Silla confirms this object as a supernova of type II, ∼ 1 month after maximum. Strong Hγ, Hβ and Hα lines with P-Cyg profiles are superimposed on a fairly blue continuum; Ti II 455.0-nm and Fe II 501.8nm and 516.9-nm lines are also present with P-Cyg profiles. The expansion velocity deduced from the Hγ and Hβ absorption is ∼ 8000 km/s. We derive a velocity for the Hα absorption of ∼ 9600 km/s. The Fe II lines show a somewhat slower expansion velocity (∼ 5800 km/s). The velocity shift of a narrow Hα emission component is consistent with the recession velocity of NGC 7650 (3300 km/s).”

1995 August 18 Brian G. Marsden